# MoS: Unleashing Parameter Efficiency of Low-Rank Adaptation with Mixture of Shards

### v1

@@ Abstract
The rapid scaling of large language models necessitates more lightweight finetuning methods to reduce the explosive GPU memory overhead when numerous customized models are served simultaneously. Targeting more parameter-efficient low-rank adaptation (LoRA), parameter sharing presents a promising solution. Empirically, our research into high-level sharing principles highlights the indispensable role of differentiation in reversing the detrimental effects of pure sharing. Guided by this finding, we propose Mixture of Shards (MoS), incorporating both inter-layer and intra-layer sharing schemes, and integrating four nearly cost-free differentiation strategies, namely subset selection, pair dissociation, vector sharding, and shard privatization. Briefly, it selects a designated number of shards from global pools with a Mixture-of-Experts (MoE)-like routing mechanism before sequentially concatenating them to low-rank matrices. Hence, it retains all the advantages of LoRA while offering enhanced parameter efficiency, and effectively circumvents the drawbacks of peer parameter-sharing methods. Our empirical experiments demonstrate approximately 8×8$\times$8 × parameter savings in a standard LoRA setting. The ablation study confirms the significance of each component. Our insights into parameter sharing and MoS method may illuminate future developments of more parameter-efficient finetuning methods.
@@ 2 Sharing & Differentiation
Reparameterization of LoRA
LoRA (Hu et al., 2021) efficiently adapts LLMs by updating the frozen pretrained weights with low-rank decomposition. Specifically, given a pretrained weight $\mathbf{W}_{0}\in\mathbb{R}^{o\times h}$, its weight update $\Delta\mathbf{W}$ is approximated as the product of two trainable matrices $\mathbf{A}\in\mathbb{R}^{r\times h}$ and $\mathbf{B}\in\mathbb{R}^{o\times r}$, where the rank $r$ is times smaller than $h$ and $o$ for parameter saving. Similar to Kopiczko et al. (2023), this product can be further reparameterized as the sum of outer products of the corresponding row and column vectors. Formally, the whole process can be presented as:
@@
Pure parameter sharing does not necessarily boost up the parameter efficiency of LoRA.
Although VeRA and PRoLoRA achieve better performance than LoRA via parameter sharing, we demonstrate that this causality is not universally applicable. To facilitate our subsequent comparisons, we introduce a straightforward sharing scheme, detailed in Sec. 3.1. Briefly, this scheme shares the low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ across all Transformer blocks for each type of linear layer. As shown in Table 1, we set the same low-level trainable parameter budget for all methods, eliminating the disturbance of parameter redundancy and enabling clear comparison of their representation capabilities. Despite several times higher rank, which is suggested for better performance (Wang et al., 2024b), this pure sharing scheme outperforms LoRA on GSM8K and Tydi QA tasks, but underperforms on others. This variability indicates that excessive sharing may hinder model performance, negating the positive effects of increased rank. In other words, sharing mechanism is not the only crucial element of existing methods for enhanced parameter efficiency.
Differentiation reverses the detrimental effects of pure sharing mechanism.
Inspired by the varying rank demands in different positions (Zhang et al., 2023), we hypothesize that the pure sharing scheme, reusing the same matrices among all Transformer blocks, contradicts different representation capacity across blocks, potentially deteriorating the performance. To verify this, we devise the following two schemes to differentiate the shared matrices:
@@
As shown in Table 1, both differentiation strategies further improve model performance over the pure sharing scheme. Specifically, random scaling can enhance the performance of pure sharing in nearly all tasks, and ultimately achieves slightly better average results. However, it still underperforms LoRA apparently. In contrast, subset selection, adopting more aggressive differential operations, reverses the degradation and consistently surpasses both pure sharing and LoRA across all benchmarks. On average, its performance exceeds that of LoRA over one percent, highlighting the importance of differentiation.
@@ 3.3 Pair Dissociation
In prior research, such as Valipour et al. (2022); Kopiczko et al. (2023), vector pairs have been regarded as the fundamental units of LoRA by default. To foster greater differentiation, we propose a vector pair disassociation strategy that is straightforward and cost-effective, requiring no additional trainable parameters while yielding significant improvements in our empirical analysis. Specifically, instead of a single shared pool for vector pairs (i.e., pairs of $\mathbf{a}_{i}^{p}$ and $\mathbf{b}_{i}^{p}$), we take decoupled vectors as the basic units, and separate $\mathbf{a}_{i}^{p}$ and $\mathbf{b}_{i}^{p}$ into two distinct pools. Hence, within the $k$-th block, the vectors that form matrices $\mathbf{A}^{k}$ and $\mathbf{B}^{k}$ will be sampled independently. Moreover, this separation naturally enables two ordered index vectors, $\mathbf{I}^{k}_{a}\in\mathbb{R}^{r}$ and $\mathbf{I}^{k}_{b}\in\mathbb{R}^{r}$, instead of unordered boolean vectors, to access vectors from the pools. Both the dissociation and the ordered indices significantly enhance the combinatorial diversity of $\mathbf{A}^{k}$ and $\mathbf{B}^{k}$. Formally, this process can be formulated as
$\Delta\mathbf{W}^{k}=\mathbf{B}^{k}\mathbf{A}^{k}=\operatorname{Index}^{c}(\mathbf{B}^{p},\mathbf{I}^{k}_{b})\operatorname{Index}^{r}(\mathbf{A}^{p},\mathbf{I}^{k}_{a}),$ (4)
where $\operatorname{Index}^{c}(\cdot)$ and $\operatorname{Index}^{r}(\cdot)$ denote the retrieval of column and row vectors, respectively, based on the given indices.
@@ 3.5 Shard Privatization
$\Delta\mathbf{W}^{k}=\mathbf{B}^{k}\mathbf{A}^{k}=\operatorname{Index}^{c}(\operatorname{Concat}(\mathbf{A}^{pub},\mathbf{A}^{pri}),\mathbf{I}^{k}_{b})\operatorname{Index}^{r}(\operatorname{Concat}(\mathbf{B}^{pub},\mathbf{B}^{pri}),\mathbf{I}^{k}_{a}),$ (5)
@@
Initialization.
Based on the shard size and the specified trainable parameter count, we first instantiate the global pools for various linear layer types, respectively. Following standard LoRA practices, we initialize $\mathbf{B}^{pub}$ and $\mathbf{B}^{pri}$ to zero, ensuring consistency with the pretrained model at the start of finetuning. As adopted in PRoLoRA (Wang et al., 2024b), we adjust the sampling boundaries of $\mathbf{A}^{pub}$ and $\mathbf{A}^{pri}$ to align with the vanilla LoRA. Subsequently, we randomly initialize the index matrices $\mathbf{I}^{k}_{a}$ and $\mathbf{I}^{k}_{b}$, and ensure the private shards are sampled only once. The training and inference procedures remain identical to those of LoRA, except that all low-rank matrices are constructed from the index matrices and global pools.
@@ 3.6 Advantage Analysis
High Parameter Efficiency.
By integrating both inter-layer and intra-layer sharing, MoS demonstrates significantly higher parameter efficiency compared to LoRA and other baselines, as shown in Sec. 4.2. This reduction in parameters results in decreased disk space and GPU memory usage, thereby greatly alleviating the serving burden in multi-LoRA scenarios.
High Representation Capacity.
As shown in Table 2, even with a rank of 256 or higher, Tied LoRA and VeRA still clearly underperform other methods. In contrast, the performance of MoS can approximately converge to that of full finetuning as do LoRA and PRoLoRA. This ensures a substantial model capacity, which is crucial for tackling challenging tasks.
@@ 4.1 General Setup
Datasets.
We finetune base models on Super-Natural Instructions (SuperNI (Wang et al., 2022)) dataset, and evaluate them on both Massive Multitask Language Understanding (MMLU (Hendrycks et al., 2021)) and TyDi QA (Clark et al., 2020) datasets for factual knowledge and multilingual capabilities, respectively. For the general and mathematical reasoning abilities, we finetune models on Flan V2 and its CoT split (Longpre et al., 2023), and conduct evaluation on Big-Bench-Hard (BBH (Suzgun et al., 2022)) and the test set of Grade School Math (GSM8K (Cobbe et al., 2021)) corpora, respectively. Moreover, coding skills are evaluated on HumanEval (Chen et al., 2021) dataset after models are finetuned on CodeAlpaca (Chaudhary, 2023) dataset.
Baselines.
From the perspective of parameter sharing, we evaluate our approach against LoRA and other existing intra-layer or inter-layer sharing methods. A brief introduction to these baselines is provided below.
@@ 4.2 Main Results
$\text{PRoLoRA}^{\dagger}$ 4/8 5.00M 45.85 36.45 27.57 49.94 36.59 19.75 36.03
MoS 4/8 5.00M 46.09 37.29 28.43 50.21 37.19 19.12 36.39
16/32 19.99M 47.01 37.79 30.93 51.71 38.34 20.00 37.63
$\text{MoS}^{\text{-sp}}$ 16/32 19.99M 46.64 36.69 30.17 50.27 36.90 18.60 36.54
$\text{MoS}^{\text{-vs}}$ 16/32 19.99M 46.47 37.52 31.77 50.90 37.98 18.69 37.22
$\text{MoS}^{\text{-pd}}$ 16/32 19.99M 46.23 36.17 30.71 51.40 37.94 16.77 36.54
@@ 4.3 Scalability Analysis
As listed in Table 3, we impose a limited trainable parameter count (equivalent to that of LoRA with the rank of 2) to compare the representational capacities of different methods. Specifically, compared to the pretrained model, LoRA achieves a notable performance improvement with an average score of 43.92. Building on this, PRoLoRA boosts the average performance by 1%. In contrast, MoS further unleashes the potential of parameters. It consistently exhibits performance enhancements across all individual benchmarks, with another increase of nearly 1% on average. This considerable improvement reinforces MoS’s positive impact on parameter efficiency and its robustness with respect to model size.
@@ 4.4 Ablation Study
Pair Dissociation.
Rather than using distinct indices for enhanced combination diversity, we apply the same index matrix for the low-rank matrix pair $\mathbf{A}^{k}$ and $\mathbf{B}^{k}$ within each linear layer (i.e., $\mathbf{I}^{k}_{a}=\mathbf{I}^{k}_{b}$). This results in an average performance decline exceeding one percent, emphasizing the critical role of pair dissociation and diverse combinations in improving model performance and aligning with our initial motivation.
Vector Sharding.
When the basic units in the global pools are reverted back to the row or column vectors of the low-rank matrices from smaller shards, only a 0.41% average decrease in performance is observed. This decline is much less than that associated with pair dissociation. Given that both components aim to enhance combination diversity, these findings imply that pair dissociation provides a more substantial boost to diversity, whereas sharding yields only marginal improvements.
Shard Privatization.
Different from the previous components, shard privatization enhances differentiation among low-rank matrices by allowing certain shards to be exclusively accessible to one low-rank matrix. Its ablation also leads to a performance drop exceeding one percent, highlighting the significance of privatization.
@@ 5 Related work
Low-rank adaptation (LoRA), introduced in Hu et al. (2021), reparameterizes weight updates with two trainable low-rank matrices while the pretrained weights remain fixed. This approach has since spurred numerous enhancements aimed at improving its effectiveness and efficiency. AdaLoRA (Zhang et al., 2023), leveraging singular value decomposition, automates rank allocation by adaptively pruning vector pairs during finetuning. However, the rank variability across layers complicates the simultaneous serving of multiple LoRAs (Sheng et al., 2023). Aimed at different parameter budget needs during inference, DyLoRA (Valipour et al., 2022) also utilizes vector pairs as fundamental units and randomly masks them throughout training, yet leading to performance degradation. Closely aligning with our work, VeRA (Kopiczko et al., 2023) implements inter-layer sharing of two frozen random matrices while updating disentangled combination vectors for each layer. However, it necessitates an excessively high rank due to limited model capacity. Subsequently, though Tied LoRA (Renduchintala et al., 2023) alleviates this constraint by allowing shared matrices to be trainable, the tying mechanism on down projection matrices restricts its applicability to linear layers of different dimensions. More recently, PRoLoRA (Wang et al., 2024b) demonstrates enhanced parameter efficiency by reusing sub-matrices within a single linear layer, thus avoiding the above limitations. However, its narrow focus on intra-layer sharing restricts its performance potential. Overall, these approaches typically assume vector pairs as the fundamental units for granted, and concentrate exclusively on inter-layer or intra-layer sharing, both of which limits the potential for parameter efficiency. In contrast, our method utilizes fine-grained shards as basic units and facilitates global sharing to achieve improved parameter efficiency.
Parameter Sharing.
@@
Parameter sharing has been extensively studied to minimize model sizes. Universal Transformer (Dehghani et al., 2018) shares all layers of a Transformer model, before Takase & Kiyono (2023) further optimizes this approach with three parameter-sharing techniques across Transformer layers for higher efficiency. DictFormer (Lou et al., 2021) implements a more compact model with a shared dictionary, unshared coefficients and indices. For on-device deployment, EdgeFormer (Ge et al., 2022) shares attention and FFN modules, and incorporates PEFT-based layer adaptation to reduce parameter count. Pires et al. (2023) eliminates the decoder layer FFN and shares a single, larger FFN across the encoder, achieving notable improvements in accuracy and latency. Recently, Cao et al. (2024) proposes two memory-efficient methods for parameter sharing in attention heads. Differently, our work is centered on the parameter sharing in LoRA modules, and explore the foundational principles guiding the development of parameter sharing strategies.
Mixture of Experts.
@@
Mixture of Experts (MoE) was initially proposed by Jacobs et al. (1991). Since Zadouri et al. (2023) replaces traditional experts with LoRA modules, various efforts have been made to perform better integration (Wu et al., 2024). Aiming at multi-task learning, Mixture-of-LoRAs (MoA) (Feng et al., 2024) integrates multiple domain-specific LoRAs through a routing strategy, improving both individual task performance and rapid domain adaptation. Similarly, MixLoRA (Li et al., 2024) constructs a sparse LoRA-based MoE model by inserting multiple LoRA experts within the FFN modules and employing independent LoRA adapters in the attention modules. Recently, Mixture of Dyadic Experts (MoDE) (Ning et al., 2024) shares a down-projection matrix for different tasks and employs atomic rank-one adapters with routers for more sophisticated task specialization, enhancing the model’s multi-tasking capabilities. In contrast, Zhou et al. (2024) present Mixture-of-Experts with Language Priors Routing (MoE-LPR), featuring a two-stage training process to improve multilingual capability without catastrophic forgetting. To the best of our knowledge, we are the first one to apply MoE for parameter savings within a single LoRA module, rather than focusing on multi-task or multilingual settings.

### v2

@@ Abstract
The rapid scaling of large language models necessitates more lightweight finetuning methods to reduce the explosive GPU memory overhead when numerous customized models are served simultaneously. Targeting more parameter-efficient low-rank adaptation (LoRA), parameter sharing presents a promising solution. Empirically, our research into high-level sharing principles highlights the indispensable role of differentiation in reversing the detrimental effects of pure sharing. Guided by this finding, we propose Mixture of Shards (MoS), incorporating both inter-layer and intra-layer sharing schemes, and integrating four nearly cost-free differentiation strategies, namely subset selection, pair dissociation, vector sharding, and shard privatization. Briefly, it selects a designated number of shards from global pools with a Mixture-of-Experts (MoE)-like routing mechanism before sequentially concatenating them to low-rank matrices. Hence, it retains all the advantages of LoRA while offering enhanced parameter efficiency, and effectively circumvents the drawbacks of peer parameter-sharing methods. Our empirical experiments demonstrate approximately 8×8$\times$8 × parameter savings in a standard LoRA setting. The ablation study confirms the significance of each component. Our insights into parameter sharing and MoS method may illuminate future developments of more parameter-efficient finetuning methods. The code is officially available at https://github.com/Forence1999/MoS.
@@ 2 Sharing & Differentiation
Reparameterization of LoRA. LoRA (Hu et al., 2021) efficiently adapts LLMs by updating the frozen pretrained weights with low-rank decomposition. Specifically, given a pretrained weight $\mathbf{W}_{0}\in\mathbb{R}^{o\times h}$, its weight update $\Delta\mathbf{W}$ is approximated as the product of two trainable matrices $\mathbf{A}\in\mathbb{R}^{r\times h}$ and $\mathbf{B}\in\mathbb{R}^{o\times r}$, where the rank $r$ is times smaller than $h$ and $o$ for parameter saving. Similar to Kopiczko et al. (2023), this product can be further reparameterized as the sum of outer products of the corresponding row and column vectors. Formally, the whole process can be presented as:
@@
Pure parameter sharing does not necessarily boost up the parameter efficiency of LoRA. Although VeRA and PRoLoRA achieve better performance than LoRA via parameter sharing, we demonstrate that this causality is not universally applicable. To facilitate our subsequent comparisons, we introduce a straightforward sharing scheme, detailed in Sec. 3.1. Briefly, this scheme shares the low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ across all Transformer blocks for each type of linear layer. As shown in Table 1, we set the same low-level trainable parameter budget for all methods, eliminating the disturbance of parameter redundancy and enabling clear comparison of their representation capabilities. Despite several times higher rank, which is suggested for better performance (Wang et al., 2024b), this pure sharing scheme outperforms LoRA on GSM8K and Tydi QA tasks, but underperforms on others. This variability indicates that excessive sharing may hinder model performance, negating the positive effects of increased rank. In other words, sharing mechanism is not the only crucial element of existing methods for enhanced parameter efficiency.
Differentiation reverses the detrimental effects of pure sharing mechanism. Inspired by the varying rank demands in different positions (Zhang et al., 2023), we hypothesize that the pure sharing scheme, reusing the same matrices among all Transformer blocks, contradicts different representation capacity across blocks, potentially deteriorating the performance. To verify this, we devise the following two schemes to differentiate the shared matrices:
@@
As shown in Table 1, both differentiation strategies further improve model performance over the pure sharing scheme. Specifically, random scaling can enhance the performance of pure sharing in nearly all tasks, and ultimately achieves slightly better average results. However, it still underperforms LoRA apparently. In contrast, subset selection, adopting more aggressive differential operations, reverses the degradation and consistently surpasses both pure sharing and LoRA across all benchmarks. On average, its performance exceeds that of LoRA over one percent, highlighting the importance of differentiation. Further experiments and analysis on LLaMA3.2-3B model (Dubey et al., 2024) are provided in Appendix B.2.
@@ 3.3 Pair Dissociation
In prior research, such as Valipour et al. (2022); Kopiczko et al. (2023), vector pairs have been regarded as the fundamental units of LoRA by default. To foster greater differentiation, we propose a vector pair disassociation strategy that is straightforward and cost-effective, requiring no additional trainable parameters while yielding significant improvements in our empirical analysis. Specifically, instead of a single shared pool for vector pairs (i.e., pairs of $\mathbf{a}_{i}^{p}$ and $\mathbf{b}_{i}^{p}$), we take decoupled vectors as the basic units, and separate $\mathbf{a}_{i}^{p}$ and $\mathbf{b}_{i}^{p}$ into two distinct pools. Hence, within the $k$-th block, the vectors that form matrices $\mathbf{A}^{k}$ and $\mathbf{B}^{k}$ will be sampled independently. Moreover, this separation naturally enables two ordered index vectors, $\mathbf{I}^{k}_{a}\in\mathbb{N}^{r}$ and $\mathbf{I}^{k}_{b}\in\mathbb{N}^{r}$, instead of unordered boolean vectors, to access vectors from the pools. Both the dissociation and the ordered indices significantly enhance the combinatorial diversity of $\mathbf{A}^{k}$ and $\mathbf{B}^{k}$. Formally, this process can be formulated as
$\Delta\mathbf{W}^{k}=\mathbf{B}^{k}\mathbf{A}^{k}=\operatorname{Route}^{c}(\mathbf{B}^{p},\mathbf{I}^{k}_{b})\operatorname{Route}^{r}(\mathbf{A}^{p},\mathbf{I}^{k}_{a}),$ (4)
where $\operatorname{Route}^{c}(\cdot)$ and $\operatorname{Route}^{r}(\cdot)$ denote the retrieval of column and row vectors, respectively, based on the given indices.
@@ 3.5 Shard Privatization
$\Delta\mathbf{W}^{k}=\mathbf{B}^{k}\mathbf{A}^{k}=\operatorname{Route}^{c}(\operatorname{Concat}(\mathbf{A}^{pub},\mathbf{A}^{pri}),\mathbf{I}^{k}_{b})\operatorname{Route}^{r}(\operatorname{Concat}(\mathbf{B}^{pub},\mathbf{B}^{pri}),\mathbf{I}^{k}_{a}),$ (5)
@@
Initialization. Based on the shard size and the specified trainable parameter count, we first instantiate the global pools for various linear layer types, respectively. Following standard LoRA practices, we initialize $\mathbf{B}^{pub}$ and $\mathbf{B}^{pri}$ to zero, ensuring consistency with the pretrained model at the start of finetuning. As adopted in PRoLoRA (Wang et al., 2024b), we adjust the sampling boundaries of $\mathbf{A}^{pub}$ and $\mathbf{A}^{pri}$ to align with the vanilla LoRA. Subsequently, we randomly initialize the index matrices $\mathbf{I}^{k}_{a}$ and $\mathbf{I}^{k}_{b}$, and ensure the private shards are sampled only once. The training and inference procedures remain identical to those of LoRA, except that all low-rank matrices are constructed from the index matrices and global pools.
Unity. Even if all the differentiation strategies are elaborated individually for clarity, they can be merged as a unified solution seamlessly for easy implementation. Briefly, MoS utilizes the routing mechanism to retrieve and concatenate shards into low-rank matrices from global pools. Additionally, we justify the theoretical motivation of each differentiation strategy from the perspective of combinational diversity in Appendix. B.1.
@@ 3.6 Advantage Analysis
High Parameter Efficiency. By integrating both inter-layer and intra-layer sharing, MoS demonstrates significantly higher parameter efficiency compared to LoRA and other baselines, as shown in Sec. 4.2. This reduction in parameters results in decreased disk space and GPU memory usage, thereby greatly alleviating the serving burden in multi-LoRA scenarios.
High Representation Capacity. As shown in Table 2, even with a rank of 256 or higher, Tied LoRA and VeRA still clearly underperform other methods. In contrast, the performance of MoS can approximately converge to that of full finetuning as do LoRA and PRoLoRA. This ensures a substantial model capacity, which is crucial for tackling challenging tasks.
@@ 4.1 General Setup
Datasets. We finetune base models on Super-Natural Instructions (SuperNI (Wang et al., 2022)) dataset, and evaluate them on both Massive Multitask Language Understanding (MMLU (Hendrycks et al., 2021)) and TyDi QA (Clark et al., 2020) datasets for factual knowledge and multilingual capabilities, respectively. For the general and mathematical reasoning abilities, we finetune models on Flan V2 and its CoT split (Longpre et al., 2023), and conduct evaluation on Big-Bench-Hard (BBH (Suzgun et al., 2022)) and the test set of Grade School Math (GSM8K (Cobbe et al., 2021)) corpora, respectively. Moreover, coding skills are evaluated on HumanEval (Chen et al., 2021) dataset after models are finetuned on CodeAlpaca (Chaudhary, 2023) dataset.
Baselines. From the perspective of parameter sharing, we evaluate our approach against LoRA and other existing intra-layer or inter-layer sharing methods. A brief introduction to these baselines is provided below.
@@ 4.2 Main Results
$\text{PRoLoRA}^{\dagger}$ 4/8 5.00M $\text{45.85}^{\ast}$ $\text{36.45}^{\ast}$ 27.57 $\text{49.94}^{\ast}$ $\text{36.59}^{\ast}$ $\text{19.75}^{\ast}$ 36.03
MoS 4/8 5.00M 46.09 37.29 28.43 $\text{50.21}^{\ast}$ $\text{37.19}^{\ast}$ 19.12 36.39
16/32 19.99M $\text{47.01}^{\ast}$ 37.79 30.93 51.71 38.34 $\text{20.00}^{\ast}$ 37.63
$\text{{MoS}{}}^{\text{-sp}}$ 16/32 19.99M $\text{46.64}^{\ast}$ 36.69 30.17 50.27 36.90 $\text{18.60}^{\ast}$ 36.54
$\text{{MoS}{}}^{\text{-vs}}$ 16/32 19.99M $\text{46.47}^{\ast}$ 37.52 31.77 50.90 37.98 $\text{18.69}^{\ast}$ 37.22
$\text{{MoS}{}}^{\text{-pd}}$ 16/32 19.99M $\text{46.23}^{\ast}$ 36.17 30.71 51.40 37.94 $\text{16.77}^{\ast}$ 36.54
@@ 4.3 Scalability Analysis
As listed in Table 3, we impose a limited trainable parameter count (equivalent to that of LoRA with the rank of 2) to compare the representational capacities of different methods. Specifically, compared to the pretrained model, LoRA achieves a notable performance improvement with an average score of 43.92. Building on this, PRoLoRA boosts the average performance by 1%. In contrast, MoS further unleashes the potential of parameters. It consistently exhibits performance enhancements across all individual benchmarks, with another increase of nearly 1% on average. Targeting smaller scale, we also extend the experiments to the LLaMA3.2-3B model with additional random seeds in Appendix. B.3, demonstrating stable performance enhancements across all individual tasks. The considerable improvement reinforces MoS’s positive impact on parameter efficiency and its robustness with respect to model sizes.
@@ 4.4 Ablation Study
Pair Dissociation. Rather than using distinct indices for enhanced combination diversity, we apply the same index matrix for the low-rank matrix pair $\mathbf{A}^{k}$ and $\mathbf{B}^{k}$ within each linear layer (i.e., $\mathbf{I}^{k}_{a}=\mathbf{I}^{k}_{b}$). This results in an average performance decline exceeding one percent, emphasizing the critical role of pair dissociation and diverse combinations in improving model performance and aligning with our initial motivation.
Vector Sharding. When the basic units in the global pools are reverted back to the row or column vectors of the low-rank matrices from smaller shards, only a 0.41% average decrease in performance is observed. This decline is much less than that associated with pair dissociation. Given that both components aim to enhance combination diversity, these findings imply that pair dissociation provides a more substantial boost to diversity, whereas sharding yields only marginal improvements.
Shard Privatization. Different from the previous components, shard privatization enhances differentiation among low-rank matrices by allowing certain shards to be exclusively accessible to one low-rank matrix. Its ablation also leads to a performance drop exceeding one percent, highlighting the significance of privatization.
@@ 5 Related work
Low-Rank Adaptation. Low-rank adaptation (LoRA), introduced in Hu et al. (2021), reparameterizes weight updates with two trainable low-rank matrices while the pretrained weights remain fixed. This approach has since spurred numerous enhancements aimed at improving its effectiveness and efficiency (Zhang et al., 2023; Wang et al., 2024a; b). AdaLoRA (Zhang et al., 2023), leveraging singular value decomposition, automates rank allocation by adaptively pruning vector pairs during finetuning. However, the rank variability across layers complicates the simultaneous serving of multiple LoRAs (Sheng et al., 2023). Closely aligning with our work, VeRA (Kopiczko et al., 2023) implements inter-layer sharing of two frozen random matrices while updating disentangled combination vectors for each layer. However, it necessitates an excessively high rank due to limited model capacity. Subsequently, though Tied LoRA (Renduchintala et al., 2023) alleviates this constraint by allowing shared matrices to be trainable, the tying mechanism on down projection matrices restricts its applicability to linear layers of different dimensions. More recently, PRoLoRA (Wang et al., 2024b) demonstrates enhanced parameter efficiency by reusing sub-matrices within a single linear layer, thus avoiding the above limitations. However, its narrow focus on intra-layer sharing restricts its performance potential. Overall, these approaches typically assume vector pairs as the fundamental units for granted, and concentrate exclusively on inter-layer or intra-layer sharing, both of which limits the potential for parameter efficiency. In contrast, our method utilizes fine-grained shards as basic units and facilitates global sharing to achieve improved parameter efficiency.
@@
Parameter Sharing. Parameter sharing has been extensively studied to minimize model sizes. Universal Transformer (Dehghani et al., 2018) shares all layers of a Transformer model, before Takase & Kiyono (2023) further optimizes this approach with three parameter-sharing techniques across Transformer layers for higher efficiency. DictFormer (Lou et al., 2021) implements a more compact model with a shared dictionary, unshared coefficients and indices. For on-device deployment, EdgeFormer (Ge et al., 2022) shares attention and FFN modules, and incorporates PEFT-based layer adaptation to reduce parameter count. Pires et al. (2023) eliminates the decoder layer FFN and shares a single, larger FFN across the encoder, achieving notable improvements in accuracy and latency. Recently, Cao et al. (2024) proposes two memory-efficient methods for parameter sharing in attention heads. Differently, our work is centered on the parameter sharing in LoRA modules, and explore the foundational principles guiding the development of parameter sharing strategies.
@@
Mixture of Experts. Mixture of Experts (MoE) was initially proposed by Jacobs et al. (1991). Since Zadouri et al. (2023) replaces traditional experts with LoRA modules, various efforts have been made to perform better integration (Wu et al., 2024). Aiming at multi-task learning, Mixture-of-LoRAs (MoA) (Feng et al., 2024) integrates multiple domain-specific LoRAs through a routing strategy, improving both individual task performance and rapid domain adaptation. Similarly, MixLoRA (Li et al., 2024) constructs a sparse LoRA-based MoE model by inserting multiple LoRA experts within the FFN modules and employing independent LoRA adapters in the attention modules. Recently, Mixture of Dyadic Experts (MoDE) (Ning et al., 2024) shares a down-projection matrix for different tasks and employs atomic rank-one adapters with routers for more sophisticated task specialization, enhancing the model’s multi-tasking capabilities. In contrast, Zhou et al. (2024) present Mixture-of-Experts with Language Priors Routing (MoE-LPR), featuring a two-stage training process to improve multilingual capability without catastrophic forgetting. To the best of our knowledge, we are the first one to apply MoE for parameter savings within a single LoRA module, rather than focusing on multi-task or multilingual settings.